what is the 90% of 80% of 75% of 600?
Answer by john500(1)

1st we take the 75%
= 600*75/100=450
2nd from the answer ie 450 we take 80 %
= 450*80/100=360
now from the second answer ie 360 we take 90%
= 360*90/100=324
So the answer is 325%